The length of a rectangle is 5 inches more than the width. The perimeter is 38 inches. Find the length and width.
Answer by tommyt3rd(5050)

1) l=w+5
2) 2(l+w)=38
from 2)
l+w=19

so that
w+5+w=19
2w=14
w=7

l=7+5=12